Kenneth Clark is 30 years and wants to retire when he is 65. So far he has saved (1) $5,750 in an IRA account in which his money is earning 8.3 percent annually and (2) $5,130 in a money market account in which he is earning 5.25 percent annually. Kenneth wants to have $1 million when he retires. Starting next year, he plans to invest the same amount of money every year until he retires in a mutual fund in which he expects to earn 7.12 percent annually. How much will Kenneth have to invest every year to achieve his savings goal?
Answer #1

Kenneth's age is 30 and he will retire at the age of 65, which means he can save the money up to 35 years.

Currently he is saving $5750 in IRA @ 8.30% annually

And $5130 in money market account @ 5.25% annually

Kenneth savings goal is 1 Million

From next year he wants to invest the same amount of money until he retires in a mutual fund where he expects to earn 7.12% annually.

IRA savings amount = $5750×8.30/100=477.25

IRA savings amount = 5750+477.25 = $6227.25

Money market account = $5130×5.25/100= 269.325

Money market savings amount = 5130+ 269.325= $5399.325

Kenneth wants to save 1 million by his retirement.

Where he already saved 6227.25+5399.325= $11626.58

Need to save after deducting the previous savings= 1000000- 11626.58 = $ 988373.4

His mutual fund interest rate is 7.12= 988373.4×7.12/100 = 70372.19

Where he has only 35 years left to earn 1 million = 70372.19/35 = $ 2010.634

He can invest $2010.634 annually to achieve his savings goal.